# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ F.2 Sensitivity Analysis
Training with 25 adv. steps 0 1 $0.16\pm 0.05$ $0.17\pm 0.06$ $0.35\pm 0.09$ $0.33\pm 0.02$ $15.26$ $0.0058$
5 $0.15\pm 0.06$ $0.16\pm 0.07$ $0.40\pm 0.09$ $0.33\pm 0.02$ $14.33$ $0.0052$
10 $0.14\pm 0.07$ $0.15\pm 0.07$ $0.43\pm 0.09$ $0.33\pm 0.02$ $14.86$ $0.0052$
20 $0.14\pm 0.06$ $0.16\pm 0.06$ $0.49\pm 0.10$ $0.32\pm 0.02$ $15.03$ $0.0047$
30 $0.12\pm 0.06$ $0.14\pm 0.05$ $0.52\pm 0.10$ $0.33\pm 0.01$ $15.46$ $0.0049$
40 $0.13\pm 0.05$ $0.14\pm 0.05$ $0.54\pm 0.10$ $0.32\pm 0.02$ $15.56$ $0.0047$
50 $0.12\pm 0.07$ $0.14\pm 0.07$ $0.54\pm 0.11$ $0.32\pm 0.02$ $15.52$ $0.0047$
Training with 25 adv. steps 50 1 $0.57\pm 0.24$ $0.59\pm 0.23$ $0.64\pm 0.26$ $0.32\pm 0.01$ $15.26$ $0.0058$
5 $0.38\pm 0.18$ $0.39\pm 0.18$ $0.51\pm 0.12$ $0.32\pm 0.01$ $14.33$ $0.0052$
10 $0.38\pm 0.16$ $0.39\pm 0.19$ $0.56\pm 0.15$ $0.32\pm 0.02$ $14.86$ $0.0052$
20 $0.39\pm 0.18$ $0.40\pm 0.18$ $0.56\pm 0.13$ $0.32\pm 0.02$ $15.03$ $0.0047$
30 $0.32\pm 0.13$ $0.36\pm 0.16$ $0.51\pm 0.12$ $0.32\pm 0.02$ $15.46$ $0.0049$
40 $0.27\pm 0.11$ $0.29\pm 0.13$ $0.53\pm 0.10$ $0.32\pm 0.02$ $15.56$ $0.0047$
50 $0.32\pm 0.15$ $0.34\pm 0.14$ $0.52\pm 0.09$ $0.33\pm 0.02$ $15.52$ $0.0047$
Training with 50 adv. steps 0 1 $0.14\pm 0.05$ $0.15\pm 0.05$ $0.33\pm 0.08$ $0.33\pm 0.02$ $15.66$ $0.0062$
5 $0.15\pm 0.07$ $0.15\pm 0.07$ $0.35\pm 0.08$ $0.33\pm 0.01$ $13.61$ $0.0047$
10 $0.13\pm 0.05$ $0.14\pm 0.06$ $0.37\pm 0.09$ $0.33\pm 0.02$ $15.16$ $0.0049$
20 $0.12\pm 0.05$ $0.13\pm 0.05$ $0.44\pm 0.08$ $0.32\pm 0.02$ $15.56$ $0.0051$
30 $0.14\pm 0.05$ $0.14\pm 0.05$ $0.45\pm 0.09$ $0.32\pm 0.02$ $15.47$ $0.0053$
40 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.50\pm 0.08$ $0.32\pm 0.01$ $16.65$ $0.0055$
50 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.52\pm 0.09$ $0.32\pm 0.01$ $16.02$ $0.0055$
Training with 50 adv. steps 50 1 $0.64\pm 0.16$ $0.69\pm 0.16$ $0.75\pm 0.20$ $0.32\pm 0.01$ $15.66$ $0.0062$
5 $0.36\pm 0.14$ $0.38\pm 0.14$ $0.54\pm 0.10$ $0.30\pm 0.02$ $13.61$ $0.0047$
10 $0.26\pm 0.15$ $0.27\pm 0.16$ $0.46\pm 0.10$ $0.30\pm 0.02$ $15.16$ $0.0049$
20 $0.29\pm 0.13$ $0.30\pm 0.13$ $0.46\pm 0.07$ $0.30\pm 0.02$ $15.56$ $0.0051$
30 $0.23\pm 0.11$ $0.28\pm 0.12$ $0.46\pm 0.10$ $0.30\pm 0.02$ $15.47$ $0.0053$
40 $0.22\pm 0.12$ $0.25\pm 0.13$ $0.48\pm 0.09$ $0.31\pm 0.02$ $16.65$ $0.0055$
50 $0.19\pm 0.10$ $0.21\pm 0.11$ $0.46\pm 0.06$ $0.31\pm 0.02$ $16.02$ $0.0055$
Training with 100 adv. steps 0 1 $0.13\pm 0.04$ $0.14\pm 0.05$ $0.29\pm 0.07$ $0.32\pm 0.02$ $15.32$ $0.0051$
5 $0.13\pm 0.05$ $0.14\pm 0.05$ $0.30\pm 0.06$ $0.32\pm 0.02$ $14.36$ $0.0049$
10 $0.13\pm 0.05$ $0.14\pm 0.06$ $0.34\pm 0.09$ $0.32\pm 0.02$ $15.56$ $0.0051$
20 $0.13\pm 0.05$ $0.13\pm 0.05$ $0.38\pm 0.08$ $0.32\pm 0.02$ $15.47$ $0.0053$
30 $0.12\pm 0.05$ $0.14\pm 0.05$ $0.42\pm 0.07$ $0.32\pm 0.02$ $15.34$ $0.0053$
40 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.43\pm 0.09$ $0.32\pm 0.02$ $16.23$ $0.0052$
50 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.46\pm 0.09$ $0.32\pm 0.02$ $17.52$ $0.0056$
Training with 100 adv. steps 50 1 $0.58\pm 0.13$ $0.61\pm 0.11$ $0.73\pm 0.13$ $0.31\pm 0.01$ $15.32$ $0.0051$
5 $0.31\pm 0.10$ $0.32\pm 0.12$ $0.44\pm 0.07$ $0.30\pm 0.02$ $14.36$ $0.0049$
10 $0.29\pm 0.11$ $0.30\pm 0.12$ $0.42\pm 0.08$ $0.29\pm 0.02$ $15.56$ $0.0051$
20 $0.25\pm 0.11$ $0.27\pm 0.11$ $0.40\pm 0.07$ $0.29\pm 0.02$ $15.47$ $0.0053$
30 $0.19\pm 0.09$ $0.20\pm 0.11$ $0.41\pm 0.08$ $0.30\pm 0.02$ $15.34$ $0.0053$
40 $0.22\pm 0.10$ $0.24\pm 0.10$ $0.41\pm 0.06$ $0.30\pm 0.02$ $16.23$ $0.0052$
50 $0.19\pm 0.10$ $0.20\pm 0.09$ $0.42\pm 0.07$ $0.30\pm 0.02$ $17.52$ $0.0056$
Setting Adv. Steps Epochs $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$ $\boldsymbol{\downarrow}\bm{\text{FID}}$ $\boldsymbol{\downarrow}\bm{\text{KID}}$
Training with 25 adv. steps 50 1 $0.55\pm 0.23$ $0.58\pm 0.22$ $0.62\pm 0.24$ $0.32\pm 0.02$ $15.26$ $0.0058$
5 $0.49\pm 0.20$ $0.49\pm 0.22$ $0.53\pm 0.16$ $0.31\pm 0.02$ $14.33$ $0.0052$
10 $0.42\pm 0.18$ $0.45\pm 0.19$ $0.54\pm 0.14$ $0.31\pm 0.02$ $14.86$ $0.0052$
20 $0.45\pm 0.18$ $0.45\pm 0.20$ $0.57\pm 0.16$ $0.31\pm 0.02$ $15.03$ $0.0047$
30 $0.35\pm 0.16$ $0.39\pm 0.17$ $0.51\pm 0.15$ $0.31\pm 0.02$ $15.46$ $0.0049$
40 $0.28\pm 0.14$ $0.29\pm 0.15$ $0.52\pm 0.12$ $0.31\pm 0.02$ $15.56$ $0.0047$
50 $0.37\pm 0.17$ $0.39\pm 0.17$ $0.52\pm 0.11$ $0.32\pm 0.02$ $15.52$ $0.0047$
Training with 25 adv. steps 100 1 $0.89\pm 0.05$ $0.88\pm 0.07$ $0.99\pm 0.01$ $0.32\pm 0.02$ $15.26$ $0.0058$
5 $0.82\pm 0.10$ $0.81\pm 0.12$ $0.89\pm 0.11$ $0.32\pm 0.02$ $14.33$ $0.0052$
10 $0.74\pm 0.15$ $0.77\pm 0.16$ $0.85\pm 0.14$ $0.32\pm 0.01$ $14.86$ $0.0052$
20 $0.76\pm 0.14$ $0.79\pm 0.14$ $0.86\pm 0.13$ $0.32\pm 0.02$ $15.03$ $0.0047$
30 $0.76\pm 0.12$ $0.77\pm 0.12$ $0.92\pm 0.08$ $0.32\pm 0.02$ $15.46$ $0.0049$
40 $0.68\pm 0.20$ $0.70\pm 0.23$ $0.63\pm 0.24$ $0.32\pm 0.02$ $15.56$ $0.0047$
50 $0.76\pm 0.12$ $0.80\pm 0.13$ $0.83\pm 0.16$ $0.32\pm 0.01$ $15.52$ $0.0047$
Training with 50 adv. steps 50 1 $0.64\pm 0.17$ $0.68\pm 0.16$ $0.72\pm 0.20$ $0.32\pm 0.01$ $15.66$ $0.0062$
5 $0.37\pm 0.12$ $0.39\pm 0.14$ $0.54\pm 0.11$ $0.30\pm 0.02$ $13.61$ $0.0047$
10 $0.26\pm 0.15$ $0.27\pm 0.16$ $0.46\pm 0.11$ $0.30\pm 0.02$ $15.16$ $0.0049$
20 $0.29\pm 0.12$ $0.30\pm 0.13$ $0.46\pm 0.07$ $0.30\pm 0.02$ $15.56$ $0.0051$
30 $0.23\pm 0.11$ $0.28\pm 0.13$ $0.46\pm 0.10$ $0.30\pm 0.02$ $15.47$ $0.0053$
40 $0.22\pm 0.12$ $0.25\pm 0.12$ $0.48\pm 0.09$ $0.31\pm 0.02$ $16.65$ $0.0055$
50 $0.19\pm 0.10$ $0.21\pm 0.11$ $0.46\pm 0.06$ $0.31\pm 0.02$ $16.02$ $0.0055$
Training with 50 adv. steps 100 1 $0.83\pm 0.08$ $0.85\pm 0.07$ $0.94\pm 0.06$ $0.32\pm 0.01$ $15.66$ $0.0062$
5 $0.54\pm 0.18$ $0.56\pm 0.17$ $0.67\pm 0.20$ $0.31\pm 0.01$ $13.61$ $0.0047$
10 $0.47\pm 0.20$ $0.45\pm 0.20$ $0.52\pm 0.17$ $0.31\pm 0.02$ $15.16$ $0.0049$
20 $0.46\pm 0.20$ $0.47\pm 0.20$ $0.61\pm 0.18$ $0.31\pm 0.02$ $15.56$ $0.0051$
30 $0.37\pm 0.19$ $0.38\pm 0.19$ $0.53\pm 0.15$ $0.31\pm 0.02$ $15.47$ $0.0053$
40 $0.40\pm 0.19$ $0.40\pm 0.21$ $0.55\pm 0.13$ $0.32\pm 0.02$ $16.65$ $0.0055$
50 $0.34\pm 0.17$ $0.34\pm 0.18$ $0.52\pm 0.10$ $0.32\pm 0.01$ $16.02$ $0.0055$
Training with 100 adv. steps 50 1 $0.59\pm 0.13$ $0.61\pm 0.10$ $0.72\pm 0.13$ $0.32\pm 0.01$ $15.32$ $0.0051$
5 $0.32\pm 0.11$ $0.33\pm 0.12$ $0.44\pm 0.07$ $0.30\pm 0.02$ $14.36$ $0.0049$
10 $0.30\pm 0.11$ $0.30\pm 0.12$ $0.42\pm 0.08$ $0.29\pm 0.02$ $15.56$ $0.0051$
20 $0.25\pm 0.11$ $0.27\pm 0.12$ $0.40\pm 0.07$ $0.29\pm 0.02$ $15.47$ $0.0053$
30 $0.19\pm 0.09$ $0.20\pm 0.11$ $0.41\pm 0.08$ $0.30\pm 0.02$ $15.34$ $0.0053$
40 $0.22\pm 0.10$ $0.24\pm 0.11$ $0.41\pm 0.06$ $0.30\pm 0.02$ $16.23$ $0.0052$
50 $0.19\pm 0.10$ $0.20\pm 0.09$ $0.42\pm 0.07$ $0.30\pm 0.02$ $17.52$ $0.0056$
Training with 100 adv. steps 100 1 $0.77\pm 0.11$ $0.79\pm 0.10$ $0.86\pm 0.13$ $0.32\pm 0.01$ $15.32$ $0.0051$
5 $0.38\pm 0.11$ $0.37\pm 0.13$ $0.48\pm 0.10$ $0.30\pm 0.02$ $14.36$ $0.0049$
10 $0.34\pm 0.13$ $0.36\pm 0.13$ $0.48\pm 0.07$ $0.30\pm 0.02$ $15.56$ $0.0051$
20 $0.26\pm 0.13$ $0.29\pm 0.14$ $0.48\pm 0.10$ $0.29\pm 0.02$ $15.47$ $0.0053$
30 $0.22\pm 0.11$ $0.23\pm 0.12$ $0.43\pm 0.08$ $0.30\pm 0.02$ $15.34$ $0.0053$
40 $0.22\pm 0.12$ $0.24\pm 0.13$ $0.48\pm 0.09$ $0.30\pm 0.02$ $16.23$ $0.0052$
50 $0.19\pm 0.10$ $0.22\pm 0.12$ $0.48\pm 0.09$ $0.31\pm 0.02$ $17.52$ $0.0056$
F.3 Ablation Study
We perform an ablation study to evaluate the impact of each individual component of the adversarial fine-tuning procedure. Specifically, we compare the default setting—where adversarial embeddings optimized for 50 steps are used over three consecutive fine-tuning steps—with alternative configurations. In one variant, only a single update is performed per embedding (One Step). In another, the model is fine-tuned exclusively on samples from the surrogate set (No Utility Loss). We also assess a setting where only non-memorized samples are used during fine-tuning (No Mitigation Loss). All configurations are evaluated across different training epochs, both when using the original training prompts without any adversarial optimization, and when using adversarial embeddings optimized for 50 steps.
The results in Tab. 11 show that fine-tuning the model for only a single step per adversarial embedding per epoch already achieves effective mitigation. This suggests that the fine-tuning process can be accelerated by reducing the number of updates per embedding. When the model is trained exclusively on surrogate samples—aiming to mitigate memorization without including any non-memorized samples to preserve utility—we observe strong mitigation, but at the cost of significantly degraded image quality, as reflected in the high FID and KID scores. Conversely, fine-tuning only on non-memorized samples while excluding surrogate samples helps maintain image quality but fails to provide sufficient mitigation against data replication.
Setting Adv. Steps Epochs $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$ $\boldsymbol{\downarrow}\bm{\text{FID}}$ $\boldsymbol{\downarrow}\bm{\text{KID}}$
Default 0 1 $0.14\pm 0.05$ $0.15\pm 0.05$ $0.33\pm 0.08$ $0.33\pm 0.02$ $15.66$ $0.0062$
5 $0.15\pm 0.07$ $0.15\pm 0.07$ $0.35\pm 0.08$ $0.33\pm 0.01$ $13.61$ $0.0047$
10 $0.13\pm 0.05$ $0.14\pm 0.06$ $0.37\pm 0.09$ $0.33\pm 0.02$ $15.16$ $0.0049$
20 $0.12\pm 0.05$ $0.13\pm 0.05$ $0.44\pm 0.08$ $0.32\pm 0.02$ $15.56$ $0.0051$
30 $0.14\pm 0.05$ $0.14\pm 0.05$ $0.45\pm 0.09$ $0.32\pm 0.02$ $15.47$ $0.0053$
40 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.50\pm 0.08$ $0.32\pm 0.01$ $16.65$ $0.0055$
50 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.52\pm 0.09$ $0.32\pm 0.01$ $16.02$ $0.0055$
Default 50 1 $0.64\pm 0.16$ $0.69\pm 0.16$ $0.75\pm 0.20$ $0.32\pm 0.01$ $15.66$ $0.0062$
5 $0.36\pm 0.14$ $0.38\pm 0.14$ $0.54\pm 0.10$ $0.30\pm 0.02$ $13.61$ $0.0047$
10 $0.26\pm 0.15$ $0.27\pm 0.16$ $0.46\pm 0.10$ $0.30\pm 0.02$ $15.16$ $0.0049$
20 $0.29\pm 0.13$ $0.30\pm 0.13$ $0.46\pm 0.07$ $0.30\pm 0.02$ $15.56$ $0.0051$
30 $0.23\pm 0.11$ $0.28\pm 0.12$ $0.46\pm 0.10$ $0.30\pm 0.02$ $15.47$ $0.0053$
40 $0.22\pm 0.12$ $0.25\pm 0.13$ $0.48\pm 0.09$ $0.31\pm 0.02$ $16.65$ $0.0055$
50 $0.19\pm 0.10$ $0.21\pm 0.11$ $0.46\pm 0.06$ $0.31\pm 0.02$ $16.02$ $0.0055$
One Step 0 1 $0.19\pm 0.05$ $0.18\pm 0.06$ $0.34\pm 0.08$ $0.33\pm 0.02$ $14.68$ $0.0056$
5 $0.13\pm 0.05$ $0.14\pm 0.05$ $0.33\pm 0.08$ $0.33\pm 0.02$ $15.07$ $0.0057$
10 $0.13\pm 0.05$ $0.14\pm 0.04$ $0.32\pm 0.07$ $0.33\pm 0.01$ $14.80$ $0.0056$
20 $0.14\pm 0.06$ $0.14\pm 0.05$ $0.37\pm 0.08$ $0.33\pm 0.02$ $14.92$ $0.0054$
30 $0.14\pm 0.05$ $0.15\pm 0.06$ $0.39\pm 0.08$ $0.33\pm 0.02$ $14.47$ $0.0045$
40 $0.12\pm 0.06$ $0.15\pm 0.06$ $0.44\pm 0.09$ $0.32\pm 0.02$ $15.75$ $0.0057$
50 $0.14\pm 0.06$ $0.14\pm 0.06$ $0.42\pm 0.07$ $0.33\pm 0.02$ $15.51$ $0.0051$
One Step 50 1 $0.69\pm 0.20$ $0.77\pm 0.16$ $0.73\pm 0.26$ $0.32\pm 0.02$ $14.68$ $0.0056$
5 $0.40\pm 0.14$ $0.43\pm 0.14$ $0.51\pm 0.13$ $0.32\pm 0.02$ $15.07$ $0.0057$
10 $0.31\pm 0.10$ $0.33\pm 0.11$ $0.47\pm 0.09$ $0.32\pm 0.02$ $14.80$ $0.0056$
20 $0.27\pm 0.10$ $0.29\pm 0.11$ $0.43\pm 0.08$ $0.32\pm 0.02$ $14.92$ $0.0054$
30 $0.26\pm 0.11$ $0.28\pm 0.12$ $0.47\pm 0.08$ $0.32\pm 0.02$ $14.47$ $0.0045$
40 $0.22\pm 0.09$ $0.23\pm 0.10$ $0.48\pm 0.09$ $0.32\pm 0.02$ $15.75$ $0.0057$
50 $0.22\pm 0.10$ $0.25\pm 0.12$ $0.49\pm 0.07$ $0.32\pm 0.02$ $15.51$ $0.0051$
No Utility Loss 0 1 $0.11\pm 0.04$ $0.13\pm 0.04$ $0.25\pm 0.08$ $0.31\pm 0.02$ $18.74$ $0.0057$
5 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.28\pm 0.08$ $0.30\pm 0.01$ $30.96$ $0.0123$
10 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.28\pm 0.06$ $0.30\pm 0.02$ $35.05$ $0.0156$
20 $0.11\pm 0.04$ $0.13\pm 0.04$ $0.38\pm 0.10$ $0.31\pm 0.02$ $63.03$ $0.0212$
30 $0.12\pm 0.05$ $0.13\pm 0.05$ $0.41\pm 0.11$ $0.30\pm 0.02$ $66.82$ $0.0237$
40 $0.10\pm 0.04$ $0.13\pm 0.04$ $0.48\pm 0.13$ $0.31\pm 0.01$ $82.38$ $0.0245$
50 $0.10\pm 0.05$ $0.12\pm 0.05$ $0.48\pm 0.10$ $0.31\pm 0.02$ $81.72$ $0.0262$
No Utility Loss 50 1 $0.42\pm 0.15$ $0.44\pm 0.16$ $0.57\pm 0.15$ $0.32\pm 0.02$ $18.74$ $0.0057$
5 $0.28\pm 0.11$ $0.30\pm 0.12$ $0.54\pm 0.08$ $0.31\pm 0.02$ $30.96$ $0.0123$
10 $0.20\pm 0.09$ $0.21\pm 0.10$ $0.52\pm 0.08$ $0.30\pm 0.02$ $35.05$ $0.0156$
20 $0.19\pm 0.08$ $0.23\pm 0.10$ $0.51\pm 0.08$ $0.30\pm 0.02$ $63.03$ $0.0212$
30 $0.20\pm 0.07$ $0.21\pm 0.08$ $0.48\pm 0.07$ $0.30\pm 0.02$ $66.82$ $0.0237$
40 $0.16\pm 0.05$ $0.19\pm 0.06$ $0.50\pm 0.08$ $0.20\pm 0.02$ $82.38$ $0.0245$
50 $0.16\pm 0.06$ $0.18\pm 0.06$ $0.47\pm 0.07$ $0.28\pm 0.03$ $81.72$ $0.0262$
No Mitigation Loss 0 1 $0.89\pm 0.06$ $0.98\pm 0.01$ $0.99\pm 0.01$ $0.33\pm 0.02$ $14.47$ $0.0056$
5 $0.86\pm 0.06$ $0.96\pm 0.01$ $0.99\pm 0.01$ $0.33\pm 0.02$ $14.45$ $0.0050$
10 $0.48\pm 0.10$ $0.57\pm 0.10$ $0.58\pm 0.13$ $0.33\pm 0.02$ $15.13$ $0.0052$
20 $0.62\pm 0.15$ $0.75\pm 0.13$ $0.64\pm 0.28$ $0.34\pm 0.02$ $14.40$ $0.0051$
30 $0.73\pm 0.11$ $0.87\pm 0.06$ $0.90\pm 0.10$ $0.33\pm 0.02$ $15.02$ $0.0046$
40 $0.63\pm 0.18$ $0.71\pm 0.19$ $0.71\pm 0.23$ $0.34\pm 0.02$ $16.04$ $0.0051$
50 $0.54\pm 0.17$ $0.65\pm 0.18$ $0.51\pm 0.11$ $0.33\pm 0.02$ $15.70$ $0.0049$
[…]

### v2

@@ I.4 Ablation Study
Default 0 1 $0.14\pm 0.05$ $0.15\pm 0.05$ $3\pm 0.08 0.33±0$$3\pm 0.02$ $0.00$ $15.66$ $0.0062$
5 $0.15\pm 0.07$ $0.15\pm 0.07$ $5\pm 0.08 0.33±0$$3\pm 0.01$ $0.00$ $13.61$ $0.0047$
10 $0.13\pm 0.05$ $0.14\pm 0.06$ $7\pm 0.09 0.33±0$$3\pm 0.02$ $0.00$ $15.16$ $0.0049$
20 $0.12\pm 0.05$ $0.13\pm 0.05$ $0.44\pm 0.08$ $32\pm 0.02 0.0$00.00 $15.56$ $0.0051$
30 $0.14\pm 0.05$ $0.14\pm 0.05$ $0.45\pm 0.09$ $32\pm 0.02 0.0$00.00 $15.47$ $0.0053$
40 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.50\pm 0.08$ $32\pm 0.01 0.0$00.00 $16.65$ $0.0055$
50 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.52\pm 0.09$ $32\pm 0.01 0.0$00.00 $16.02$ $0.0055$
Default 50 1 $0.64\pm 0.16$ $0.69\pm 0.16$ $0.75\pm 0.20$ $0.32\pm 0.01$ $0.17$ $15.66$ $0.0062$
5 $0.36\pm 0.14$ $0.38\pm 0.14$ $0.54\pm 0.10$ $30\pm 0.02 0.0$20.02 $13.61$ $0.0047$
10 $0.26\pm 0.15$ $0.27\pm 0.16$ $0.46\pm 0.10$ $30\pm 0.02 0.0$10.01 $15.16$ $0.0049$
20 $0.29\pm 0.13$ $0.30\pm 0.13$ $0.46\pm 0.07$ $30\pm 0.02 0.0$00.00 $15.56$ $0.0051$
30 $0.23\pm 0.11$ $0.28\pm 0.12$ $0.46\pm 0.10$ $30\pm 0.02 0.0$00.00 $15.47$ $0.0053$
40 $0.22\pm 0.12$ $0.25\pm 0.13$ $0.48\pm 0.09$ $31\pm 0.02 0.0$00.00 $16.65$ $0.0055$
50 $0.19\pm 0.10$ $0.21\pm 0.11$ $0.46\pm 0.06$ $31\pm 0.02 0.0$00.00 $16.02$ $0.0055$
One Step 0 1 $0.19\pm 0.05$ $0.18\pm 0.06$ $4\pm 0.08 0.33±0$$3\pm 0.02$ $0.00$ $14.68$ $0.0056$
5 $0.13\pm 0.05$ $0.14\pm 0.05$ $3\pm 0.08 0.33±0$$3\pm 0.02$ $0.00$ $15.07$ $0.0057$
10 $0.13\pm 0.05$ $0.14\pm 0.04$ $2\pm 0.07 0.33±0$$3\pm 0.01$ $0.00$ $14.80$ $0.0056$
20 $0.14\pm 0.06$ $0.14\pm 0.05$ $7\pm 0.08 0.33±0$$3\pm 0.02$ $0.00$ $14.92$ $0.0054$
30 $0.14\pm 0.05$ $0.15\pm 0.06$ $9\pm 0.08 0.33±0$$3\pm 0.02$ $0.00$ $14.47$ $0.0045$
40 $0.12\pm 0.06$ $0.15\pm 0.06$ $0.44\pm 0.09$ $32\pm 0.02 0.0$00.00 $15.75$ $0.0057$
50 $0.14\pm 0.06$ $0.14\pm 0.06$ $0.42\pm 0.07$ $33\pm 0.02 0.0$00.00 $15.51$ $0.0051$
One Step 50 1 $0.69\pm 0.20$ $0.77\pm 0.16$ $0.73\pm 0.26$ $0.32\pm 0.02$ $0.49$ $14.68$ $0.0056$
5 $0.40\pm 0.14$ $0.43\pm 0.14$ $0.51\pm 0.13$ $0.32\pm 0.02$ $0.10$ $15.07$ $0.0057$
10 $\pm 0.10 0.33±0.11$$3\pm 0$.11 $0.47\pm 0.09$ $32\pm 0.02 0.0$30.03 $14.80$ $0.0056$
20 $0.27\pm 0.10$ $0.29\pm 0.11$ $0.43\pm 0.08$ $32\pm 0.02 0.0$20.02 $14.92$ $0.0054$
30 $0.26\pm 0.11$ $0.28\pm 0.12$ $0.47\pm 0.08$ $32\pm 0.02 0.0$10.01 $14.47$ $0.0045$
40 $0.22\pm 0.09$ $0.23\pm 0.10$ $0.48\pm 0.09$ $32\pm 0.02 0.0$20.02 $15.75$ $0.0057$
50 $0.22\pm 0.10$ $0.25\pm 0.12$ $0.49\pm 0.07$ $32\pm 0.02 0.0$00.00 $15.51$ $0.0051$
No Utility Loss 0 1 $0.11\pm 0.04$ $0.13\pm 0.04$ $0.25\pm 0.08$ $31\pm 0.02 0.0$00.00 $18.74$ $0.0057$
5 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.28\pm 0.08$ $30\pm 0.01 0.0$00.00 $30.96$ $0.0123$
10 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.28\pm 0.06$ $30\pm 0.02 0.0$00.00 $35.05$ $0.0156$
20 $0.11\pm 0.04$ $0.13\pm 0.04$ $0.38\pm 0.10$ $31\pm 0.02 0.0$00.00 $63.03$ $0.0212$
30 $0.12\pm 0.05$ $0.13\pm 0.05$ $0.41\pm 0.11$ $30\pm 0.02 0.0$00.00 $66.82$ $0.0237$
40 $0.10\pm 0.04$ $0.13\pm 0.04$ $0.48\pm 0.13$ $31\pm 0.01 0.0$00.00 $82.38$ $0.0245$
50 $0.10\pm 0.05$ $0.12\pm 0.05$ $0.48\pm 0.10$ $31\pm 0.02 0.0$00.00 $81.72$ $0.0262$
No Utility Loss 50 1 $2\pm 0.15 0.44±0$$4\pm 0.16$ $0.57\pm 0.15$ $32\pm 0.02 0.0$90.09 $18.74$ $0.0057$
5 $0.28\pm 0.11$ $0.30\pm 0.12$ $0.54\pm 0.08$ $31\pm 0.02 0.0$00.00 $30.96$ $0.0123$
10 $0.20\pm 0.09$ $0.21\pm 0.10$ $0.52\pm 0.08$ $30\pm 0.02 0.0$00.00 $35.05$ $0.0156$
20 $0.19\pm 0.08$ $0.23\pm 0.10$ $0.51\pm 0.08$ $30\pm 0.02 0.0$00.00 $63.03$ $0.0212$
30 $0.20\pm 0.07$ $0.21\pm 0.08$ $0.48\pm 0.07$ $30\pm 0.02 0.0$00.00 $66.82$ $0.0237$
40 $0.16\pm 0.05$ $0.19\pm 0.06$ $0.50\pm 0.08$ $20\pm 0.02 0.0$00.00 $82.38$ $0.0245$
50 $0.16\pm 0.06$ $0.18\pm 0.06$ $0.47\pm 0.07$ $28\pm 0.03 0.0$00.00 $81.72$ $0.0262$
No Mitigation Loss 0 1 $0.89\pm 0.06$ $8\pm 0.01 0.99±0$$9\pm 0.01$ $0.33\pm 0.02$ $0.91$ $14.47$ $0.0056$
5 $0.86\pm 0.06$ $6\pm 0.01 0.99±0$$9\pm 0.01$ $0.33\pm 0.02$ $0.86$ $14.45$ $0.0050$
10 $0.48\pm 0.10$ $0.57\pm 0.10$ $0.58\pm 0.13$ $33\pm 0.02 0.0$60.06 $15.13$ $0.0052$
20 $0.62\pm 0.15$ $0.75\pm 0.13$ $0.64\pm 0.28$ $0.34\pm 0.02$ $0.30$ $14.40$ $0.0051$
30 $0.73\pm 0.11$ $0.87\pm 0.06$ $0.90\pm 0.10$ $0.33\pm 0.02$ $0.55$ $15.02$ $0.0046$
40 $0.63\pm 0.18$ $0.71\pm 0.19$ $0.71\pm 0.23$ $\pm 0.02 0.340.34$ $16.04$ $0.0051$
50 $0.54\pm 0.17$ $0.65\pm 0.18$ $0.51\pm 0.11$ $0.33\pm 0.02$ $0.21$ $15.70$ $0.0049$
No Mitigation Loss 50 1 $0.91\pm 0.03$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $1.00$ $14.47$ $0.0056$
5 $0.90\pm 0.03$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $0.98$ $14.45$ $0.0050$
10 $0.88\pm 0.03$ $0.92\pm 0.04$ $1.00\pm 0.00$ $0.32\pm 0.02$ $0.96$ $15.13$ $0.0052$
20 $0.88\pm 0.04$ $0.94\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.02$ $0.98$ $14.40$ $0.0051$
30 $0.88\pm 0.04$ $0.94\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.01$ $0.97$ $15.02$ $0.0046$
40 $0.85\pm 0.04$ $0.92\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.01$ $0.91$ $16.04$ $0.0051$
50 $0.86\pm 0.05$ $0.92\pm 0.04$ $1.00\pm 0.00$ $0.32\pm 0.02$ $0.91$ $15.70$ $0.0049$
I.5 Cost Discussion
Our mitigation method can seem costly at first, given it requires generating a set of surrogate images from the model, and performing a full fine-tuning of the whole DM. In total, the runtime of our method on a single A100 GPU can be approximated with the following equation
$\text{Time}=(\text{N}\times 20)+(\text{E}\times\text{N}\times 10)\text{,}$ (5)
where Time is in seconds, N is the number of memorized images to remove, and E is the number of epochs we run our method. For N=112, E=5, we get about 2.5h of runtime, which we consider reasonable, given the effectiveness of our method.
An alternative approach would be to apply LoRA (Hu et al., 2021) adaptation on the DM, to lower the compute cost. We use LoRA with rank 128 on (1) all fully-connected layers, and (2) both fully-connected and convolutional layers. Unfortunately, in neither of those cases the memorization is fully mitigated, even after 50 epochs of our mitigation method. We are still able to craft adversarial embeddings to replicate memorized data. This result is consistent with our earlier findings: memorization is not confined to individual weights or layers, but is distributed across the entire model and requires updates to additional components.
I.6 Our Mitigation is Robust to Other Adversarial Optimization Techniques
While our adversarial fine-tuning is robust against Dori, we acknowledge that it might not be robust to other methods of obtaining adversarial embeddings. Since Dori is the first such method for triggering memorization, we apply UnlearnDiffAtk (Zhang et al., 2024c) against our fine-tuned model. See Section˜G.1 for details on the implementation. The prompt optimization algorithm is GCG (Zou et al., 2023), arguably the strongest adversarial optimization method available at the time of writing the paper. We find that our model is robust against UnlearnDiffAtk, i.e., the adversarial prompts do not cause replication of memorized content.
[…]